# Rui Eduardo Lopes, Student Member, IEEE, Susana Sargento, Senior Member, IEEE

### v1

@@ V-A Towards autonomous driving for all road situations
Context Aware Maneuvering Vehicles may consider cooperative manoeuvres between non-autonomous vehicles and/or autonomous vehicles to ensure complex manoeuvres are done with no harm.
For example, within the context of the cooperation between public and private transportation, a service can be considered to give priority to bus, ambulance or other public transportation to enter a certain crossroad instead of a privately owned vehicle. This service shall be dynamic and consider several potential alternatives, including scenarios where the proprietary vehicle must in fact stop in order to prevent an accident. The service(s) to decide each vehicle’s action require a minimized bounded latency to receive information on the road situation and to send information with the proposed action to other vehicle(s).
When considering the air mobility paradigm, in particular electrical Vertical Take Off and Landing (eVTOL) vehicles, the cooperation between vehicles gain an additional importance. New scenarios shall be considered, including cooperation between vehicles and between the eVTOL and the road infrastructure — such as landing pads [9] — to ensure no accident occurs during take off or landing, even in bad visibility (by occlusion or lack of lighting). In these scenarios coordination between vehicles and infrastructure can avoid a catastrophic air accident between the vehicles or a crash on top of property or people.
@@
Autoware, an open source framework considered to be state-of-the-art for autonomous driving [^54] Apart from closed source solutions from automakers (more details on section II-C), where it is not possible to fully understand the scope of the cooperation between vehicles and its interoperability with other brands., considers several services to ensure an autonomous driving — albeit in previously mapped environments — to support manoeuvres such as curb-to-curb, lane following, lane merging, parking or reaction to traffic lights and dynamic entities (e.g. Vulnerable Road Users).
@@
To produce plans that eventually become decisions and actuation to the engines, braking and other vehicular subsystems, Autoware relies on information of the vehicle itself and on a model of the environment. All this planning is therefore very local, with cooperation from other vehicles and road entities proving essential not only to improve Autoware’s vision of the world (in a cooperative sensing solution such as the one defined by ETSI Collective Perception Service (CPS)), but also to coordinate the manoeuvre that each vehicle will do, ensuring that a safe manoeuvre for one vehicle does not possess danger to another.
@@
When the planning becomes cooperative and not only local, a new problem arises. Information from and to each vehicle must be sent and must be sent within a well defined time-frame that is compatible with Autoware producing a decision and actuating on the vehicle’s subsystems. If the information arrives after this defined time-frame, the decision will not be correct, while if it arrives before this defined time-frame, the decision might be unnecessary — but also potentially dangerous [^55] For example, a vehicle would end up braking before being necessary, becoming vulnerable to being hit in the back..
This rational defines the currently considered use case: cooperative deterministic manoeuvreing, where vehicles receive information from other entities and use them as part of the planning.
@@
As an example scenario, one may consider coordinating lane merging manoeuvre between an ambulance and a private vehicle. This use case requires the development of services depicted in fig. 15, namely a) a message creator and producer to support exchanging data between Autoware instances in different vehicles, and b) a parser and injector of such information into Autoware planning. The data format would follow ETSI MCM standards and it would be periodic, with deadline insurance requirements — and therefore both a minimum and bounded latency requirement — with no tolerance to failures of delivery and with a variable size with an upper bounds corresponding to the maximum size of the MCM message.
@@
The integration of Autoware with cooperative manoeuvreing is considered in [138] [^56] Additional details at https://tlab-wide.github.io/AutowareV2X/pr-48/design/use-cases/.. However, it fails to consider the deterministic communications requirements or the need for orchestrating the overall deployment in vehicles’ ECUs in a more automated approach.
@@
Vehicle malfunction situations Even when full autonomy of a vehicle is not only a reality but also legal, vehicles can still have unexpected issues that endanger their occupants and surroundings. Examples include a flat tire, a tire explosion, failed waterproofing of the batteries (in the case of electric vehicles), or a failure on the valve timing (in the case of internal combustion engines). In these scenarios, services to analyze, monitor and warn others about the situation should be deployed in a bounded minimal time and the information to other vehicles shall be sent with both minimized, and bounded latency.
@@
When considering Connected Autonomous Vehicles, services that monitor and react when sensors are malfunctioning are critical and their information must also be bounded in time and transmitted reliably. Without such services, the CAVs cannot react to a malfunctioning or occluded camera, LIght Detection And Ranging (Lidar) or equivalent and can wrongly assume their information is correct, with potentially deadly consequences.
V-B Towards vehicles tailored for everyone
@@
While the future of mobility is inherently more autonomous and more feature-packed, less concern is given to occupants, including children or occupants who may have some sort of physical or cognitive restriction. Restrictions considered here may include some minor impairment, vision or hearing issues, concentration issues, and others, that restrict the occupants’ capability of reacting to the environment in the same way.
@@
The following situations are examples where a vehicle has an occupant with some issue/limitation who can greatly benefit from a SDV paradigm. In this paradigm, the vehicle can deploy services tailored for the people’s specific needs for customization of vehicular functions, protection, monitoring, and companionship. In all situations, the services to be deployed shall have their data delivered in a bounded time — so the services are not left in an inconsistent state.
@@
Automatic vehicle configuration based on occupants characteristics Information on the occupants’ characteristics and vital parameters can be obtained by intra-vehicular subsystems that allow the vehicle to adapt interior functions (e.g. lighting, sound, infotainment) to the best comfort of the occupants, therefore creating a set of customizations tailored to each occupant’s characteristics.
@@
Not only the services can adapt but warn if there is an issue (like a non-fixed wheelchair or baby seat). Such warnings can be sent both to the vehicle’s infotainment systems but also to other vehicles. Vehicles can then use such information, as well as information about the occupants’ characteristics, to deploy services to adapt the vehicle driving — e.g. to increase cruise control following distances, reduce throttle response, or warn other smart city infrastructures.
@@
These services can greatly benefit the healthcare industry in particular. Consider a scenario where an ambulance is transporting a patient. Vehicular services can support scenarios such as this:k
@@
  • Other vehicles: OK, I need to slow down.
  • Cloud: I’ll provide you with specific applications (e.g. to control acceleration profiles or climate control settings). I’ll deploy an application for monitoring impaired people that will run on a cluster of $X$ vehicles.
Adapting the vehicle’s subsystems to impaired occupants Consider a situation where, in order to transport impaired people in a wheelchair, a healthcare provider must [^57] From an informal interview with a specialist nurse in medical-surgical nursing:
@@
    Get the wheel chair up a moving ramp or lift, per depicted in footnote 59585858Another example of such ramps and lifts in https://www.gilaniengineering.com.au/product/hydraulic-wheelchair-platform-accessible-lift-hoist-for-cars-and-vans/.
@@
    Attach the wheelchair to the vehicle through a set of several belts fixed to specific points within the vehicle, as depicted in footnote 60.
(a) Example of ramp where wheelchairs need to be pushed manually by the healthcare provider (courtesy of Tripod Mobility [^59] More details at https://www.tripodmobility.com/en/products/wheelchair-accesible-vehicles/upfront/.).
(b) Example of expected attachment points within a vehicle for a wheelchair (courtesy of BraunAbility [^60] Available at https://www.braunability.eu/en/products/tie-downs-and-seatbelts/wtors/.).
@@
The healthcare provider must do such manually, multiple times (for multiple people) and in a short time-frame and without little to no feedback if the person and the chair is actually fixed properly. Mistakes include fixing outside of the fixing point, twisting the straps and/or simply not fixing the person and/or chair. In addition to that, specific formation is required to do such tasks.
@@
A vehicular services (or set of services) that is capable of communicating with other in vehicle systems (e.g. infotainment), but also with the healthcare provider registry systems, would greatly decrease the potential human errors of the current solution, and since it would minimize the dependency on humans could even support an increase on the patient’s privacy. One could even consider a more intelligent connected wheelchair directly coordinating with the vehicle all the necessary movements. Such systems would not require a minimized latency but guarantees of delivery of the information within well-known time bounds and with low jitter in order to ensure no unwanted movements of the wheelchair that may harm the person.
Other situations include:
@@
  • Monitoring a baby that is ill or very restless and provide such information directly to a hospital.
  • Decrease the brightness of other vehicle’s lightning in order to protect the eyes of a newly operated occupant.
@@
Personal assistants While the SDV paradigm is parallel to the advent of autonomous vehicles, vehicles can include humans in new ways within the driving experience. The unparalleled amount of information and sensing the vehicle can obtain from its environment, driver and occupants, can be used to support personal assistants that give occupants the sensation of being in control — even if the vehicle is self-driving.
@@
Such assistants can be developed as a set of subscribing services within the vehicle’s ECUs dedicated to infotainment subsystems and, if using speech or touch interaction modalities, can greatly benefit from a well defined timely delivery of information which decreases the overall jitter visible to the end users.
V-C Towards assisted fleet management and maintenance
@@
Remote vehicle health and predictive maintenance Constant monitoring of the main vital parameters of a vehicle state — such as tire condition, brake system condition, engine cooling system status (in the case of internal combustion engines) or the power delivery subsystems (in the case of electric vehicles) — requires not only a reduced latency but also, in case the vehicle detects an issue, a guarantee of delivery in a bounded time so at the very least other vehicles are aware of the problem before it has impact on the road. In the case of emergency scenarios — as previously described in fig. 15, the latency shall also be minimized.
@@
Fleet management Fleet management scenarios and remote services, such as the one presented at fig. 17616161More details at https://www.autowiz.in/fleet-features.html., are other examples of use cases that can benefit from a set of services deployed across multiple vehicles. Considering a set of services containing the logic required to monitor, in real-time, the status (e.g. location or number of hours the vehicle is driven) of a vehicle within a fleet. This set of services, if orchestrated throughout a cluster of vehicles, can support systems for enterprises managing a vehicular fleet, homologating it, or even for state entities to perform the routine inspection. Information from/to those services, while do not necessarily require an extremely low latency, must be delivered within known time bounds and with low jitter.
@@
Over-the-air updates and real-time homologation Changes in the vehicles’ software — either via OTA updates, through the addition of new services or after a vehicle is sold to ensure no changes — should, in an automated way, be reflected in a re-homologation process. This process, while not requiring a minimized latency, shall be done in a bounded timely manner, with a guarantee of reliability, otherwise, vehicles may be driving on public roads with an unwanted, unsafe state [5].
VI Proposed vision
@@
In order to support a dynamic management of the services several layers must be considered as depicted in fig. 18. These layers should act at least in the following domains [32]:
@@
  • Orchestration of the services: where they should be run and policies to decide the node(s) and the number of replicas.
@@
  • Virtualization and management of the network: including automatic configuration of routing and TSN shapers and reliability processes (e.g. FRER).
@@
  • Virtualization of the hardware in order to support better resource usage and decoupling of services and operating systems from each ECU.
@@
Figure 18: Overall layered view of the architecture main modules.
@@
Figure 19 presents the overview of an initial proposal of architecture. This architecture supports a cloud computing approach on vehicles to see them as a data center on wheels [32] in order to support service scheduling, redundancy, and healing based on the vehicles’ computing capabilities.
@@
Figure 19: Overall high level architecture of the proposed solution.
@@
The Orchestrator vertical (in yellow) is the main entry-point of the system to be researched and ensures the first aforementioned layer. Through user-defined and system-defined metrics within a Service Descriptor, decides the amount of resources and the placement of those resources throughout the computing continuum and the different vehicles throughout the smart city, while considering possible handovers and optimal usage of the computing and network resources. The orchestrator is responsible for controlling 4 other verticals.
@@
The first vertical, the Deterministic Network (in blue), ensures that messages – either control messages to actuate or status messages containing information from vehicles and the smart city – are delivered in a well-specified time bound and with fault-proof reliability. Research topics here include, but are not limited to, integrating clock synchronization mechanisms into wireless vehicular networks and mechanisms towards dynamic automatic definition, and optimal configuration of flows of packets, their queue priorities and characteristics. The module within the architecture, the Deterministic Network Configurator, is responsible for creating the necessary traffic flows, their parameters and configuring the necessary hardware, namely the Network Interface Controller/Card (NIC) within each ECU and the TSN capable Ethernet switches interconnecting them. It is expected at least 3 traffic classes to be configured: a) system control messages with maximum criticality; b) high size (non) periodic streams of data and c) periodic or non periodic service messages.
@@
The Data Layer vertical (in green) is responsible for ensuring the data is exchanged between the vehicular services in the most appropriated set of SOA protocols — such as SOME/IP, MQTT or DDS.
@@
The Hypervisor vertical (in red) is responsible for ensuring the deployment of the services, either within a Kubernetes worker node, within a virtual machine managed for example by KVM or through a containerd or equivalent container runtime.
@@
These 3 verticals dispatch orders to the Vehicle Abstraction Layer (VAL) (in purple) of each ECU, which abstracts different types and models of vehicles’ ECUs – both simulators and real vehicles – their actuators, sensors, and ECUs through the abstraction of existing frameworks, such as AUTOSAR, Autoware or others. Conceptually, the VAL is no different to the standardized abstractions operating systems created during the mid 1980s and 1990s to abstract the hardware from the remaining kernel, with [10] proposing an extension of such a paradigm into the vehicular paradigm, as depicted in fig. 20.
Figure 20: The VAL based architecture for vehicles as described by [10].
@@
VI-A Overall end-to-end interaction
Figure 21 provides an overview of the end-to-end interaction of the system proposal, from the features the end-user (developer of vehicular services) can expect to see up to the configuration of the necessary hardware for the services to be run while fulfilling their requirements.

### v2

@@ 5.2 Services for autonomous driving in all road situations
5.2.1 Context Aware Maneuvering
@@
Vehicles may consider cooperative manoeuvres between non-autonomous vehicles and/or autonomous vehicles to ensure that complex manoeuvres are done with no harm. For example, within the context of the cooperation between a personal vehicle public and private transportation, a service can be considered to give priority to bus, ambulance or other public transportation to enter a certain crossroad instead of a privately owned vehicle. This service shall be react to changes on the environment and surroundings of the vehicle and decide the action of each vehicle. Both receive information on the road situation and sending the proposed action to other vehicles would require a minimized bounded latency to ensure a proper decision and that such decision is enforced in the right moment.
When considering the air mobility paradigm, in particular electrical Vertical Take Off and Landing (eVTOL) vehicles, the cooperation between the vehicle and other vehicles gain an additional importance. New scenarios shall be considered, including cooperation between vehicles and the eVTOL and the road infrastructure (such as landing pads) to ensure that no accident occurs during take off or landing, even in bad visibility (by occlusion or lack of lighting). In these scenarios, coordination between vehicles and infrastructure can avoid a catastrophic air accident between the vehicles or a potentially deadly crash on top of property or people [7].
@@
Autoware, an open source framework considered to be state-of-the-art for autonomous driving [^53] Apart from closed source solutions from automakers (more details on section 2.1), where it is not possible to fully understand the scope of the cooperation between vehicles and its interoperability with other brands., considers several services to ensure an autonomous driving — albeit in previously mapped environments — to support manoeuvres such as curb-to-curb, lane following, lane merging, parking or reaction to traffic lights and dynamic entities (e.g. Vulnerable Road Users).
@@
To produce plans that eventually become decisions and actuation to the engines, braking and other vehicular subsystems, Autoware relies on information of the vehicle itself and on a model of the environment. All this planning is therefore very local, with cooperation from other vehicles and road entities proving essential not only to improve Autoware’s vision of the world —in a cooperative sensing solution such as the one defined by ETSI Collective Perception Service (CPS)—, but also to coordinate the manoeuvre that each vehicle will do, ensuring that a safe manoeuvre for one vehicle does not possess danger to another.
@@
When the planning becomes cooperative and not only local, a new problem arises. Information from and to each vehicle must be sent within a well defined time-frame that is compatible with Autoware producing a decision and actuating on the vehicle’s subsystems. If the information arrives after this defined time-frame, the decision will not be correct, while if it arrives before this defined time-frame, the decision might be unnecessary — but also potentially dangerous [^54] For example, a vehicle would end up braking before being necessary, becoming vulnerable to being hit in the back..
@@
As an example scenario, one may consider coordinating lane merging manoeuvre between an ambulance and a private vehicle. This use case requires the development of services depicted in fig. 17, namely a) a message creator and producer to support exchanging data between Autoware instances in different vehicles, and b) a parser and injector of such information into Autoware planning. The data format would follow ETSI MCM standards and it would be periodic, with deadline insurance requirements (and therefore both a minimum and bounded latency requirement) with no tolerance to failures of delivery, and with a variable size with an upper bounds corresponding to the maximum size of the MCM message.
@@
The integration of Autoware with cooperative manoeuvering is considered in [139] [^55] Additional details at tlab-wide.github.io/AutowareV2X.. However, it fails to consider the deterministic communications requirements or the need for orchestrating the overall deployment in vehicles’ ECUs in a more automated approach.
5.2.2 Vehicle malfunction situations
@@
Even when full autonomy of a vehicle is not only a reality but also legal, vehicles can still have unexpected issues that endanger their occupants and surroundings. Examples include a flat tire, a tire explosion, failed waterproofing of the batteries (in the case of electric vehicles), or a failure on the valve timing (in the case of internal combustion engines). In these scenarios, services to analyze, monitor and warn others about the situation should be deployed in a bounded minimal time, and the information to other vehicles shall be sent with both minimized, and bounded latency.
@@
When considering Connected Autonomous Vehicles, services that monitor and react when sensors are malfunctioning are critical, and their information must also be bounded in time and transmitted reliably. Without such services, the CAVs cannot react to a malfunctioning or occluded sensor (such as camera, LIght Detection And Ranging (Lidar) or equivalent) and can wrongly assume that their information is correct, with potentially deadly consequences.
5.3 Services for vehicles tailored for everyone
@@
While the future of mobility is seemingly more autonomous and more feature-packed, less concern is given to occupants and on their needs, including children or occupants who may have some sort of physical or cognitive restriction. Restrictions considered here may include some minor impairment, vision or hearing issues, concentration issues, and others, that restrict the occupants’ capability of reacting to the environment in the same way.
@@
The following situations are examples where a vehicle has an occupant with some issue/limitation who can greatly benefit from a SDV paradigm. In this paradigm, the vehicle can deploy services tailored for the people’s specific needs for customization of vehicular functions, protection, monitoring, and companionship. In all situations, like in previous use cases, the services to be developed and deployed shall have their data delivered in a bounded time — so the services are not left in an inconsistent state.
5.3.1 Automatic vehicle configuration based on occupants characteristics
@@
Information on the occupants’ characteristics and vital parameters can be obtained by intra-vehicular subsystems that allow the vehicle to adapt interior functions (e.g. seating position and configurations, lighting, sound, infotainment) to the best comfort of the occupants, therefore creating a set of customizations tailored to each occupant’s characteristics.
@@
Not only the services can adapt, but warn if there is an issue (like a non-fixed wheelchair or baby seat). Such warnings can be sent both to the infotainment vehicular subsystems, and also to other vehicles. Vehicles can then use such information, as well as information about the occupants’ characteristics, to deploy services to adapt the vehicle driving mode, profile and characteristics. For example, such adaptation could signify an increase cruise control following distances, reduce throttle response, or warn other smart city infrastructures when someone inside the vehicle is ill or impaired.
@@
These services can greatly benefit the healthcare industry in particular. Consider a scenario where an ambulance is transporting a patient. Vehicular services can support scenarios such as this:
  • 1.
@@
  • 2.
    Other vehicles: Understood, I need to slow down and provide space around the ambulance.
  • 3.
    SDV Orchestrator: We’ll provide you with specific applications (e.g. to control acceleration profiles or climate control settings). We’ll deploy an application for monitoring ill people that will run on a cluster of $X$ vehicles. We will coordinate with all the vehicles in the path, and with the city traffic manager, to create a corridor for you to be able to pass between other vehicles in the city.
5.3.2 Adapting the vehicle’s subsystems to impaired occupants
Consider a situation where, in order to transport impaired people in a wheelchair, a healthcare provider must [^56] From an informal interview with a specialist nurse in medical-surgical nursing.:
@@ 5.3.2 Adapting the vehicle’s subsystems to impaired occupants
    Get the wheel chair up a moving ramp or lift, per depicted in footnote 58575757Another example of such ramps and lifts in gilaniengineering.com.au..
@@
    Attach the wheelchair to the vehicle through a set of several belts fixed to specific points within the vehicle, as depicted in footnote 59.
(a) Example of ramp where wheelchairs need to be pushed manually by the healthcare provider (courtesy of Tripod Mobility [^58] More details at tripodmobility.com.).
(b) Example of expected attachment points (4 in this case) within a vehicle for a wheelchair (courtesy of BraunAbility [^59] Available at braunability.eu.).
@@
The healthcare provider must do such manually, multiple times (for multiple people) and in a short time-frame and without little to no feedback on whether the person and the chair are actually fixed properly. Mistakes include fixing outside of the fixing point(s), twisting the straps and/or simply not fixing the person and/or chair. In addition to that, specific formation is required to do such tasks.
@@
A vehicular service (or set of services) that is capable of communicating with other in-vehicle systems (e.g. infotainment), but also with the healthcare provider registry systems, would greatly decrease the potential human errors of the current solution, and since it would minimize the dependency on humans, it could even support an increase on the patient’s privacy. One could even consider a more intelligent connected wheelchair directly coordinating with the vehicle all the necessary movements. Such systems would not require a minimized latency, but guarantees of delivery of the information within well-known time bounds and with low jitter in order to ensure no unwanted movements of the wheelchair that may harm the person.
The adaptability of the vehicle does not need to end on improving or automate mechanical subsystems. In fact, greater challenges can be considered. Depending on the profile of the person (whether it’s any sort of impairment, special needs or old age), the vehicle could further adapt itself for example by choosing a softer, appropriated music (supporting music theraphy [^60] More details at britannica.com/topic/music-therapy.), adapt the climate subsystems to ensure the adequate temperature, tint the glass to avoid too much sun exposure.
Other scenarios may include the protection of other type of vulnerable humans, specifically babies:
  • 1.
    Managing situations were babies are left unattended in cars. It is a widely known issue [^61] Unfortunate examples at cnn.com.. Babies, when left in vehicles and especially in hot weather, where vehicle temperatures can rise rapidly, are at risk of heatstroke and potential death. Vehicular subsystems and services could detect the presence of the baby and alert caregivers or authorities if the child is left inside under dangerous conditions.
  • 2.
    Managing the set of airbags that should or should not be deployed based on the position or age of the baby.
  • 3.
@@
    Monitoring a baby who is ill or very restless and provide such information directly to a hospital.
5.3.3 Personal assistants
@@
Vehicles can include humans in new ways within the driving experience, even when considering the SDV paradigm is parallel to the advent of autonomous vehicles. The unparallel amount of information and sensing the vehicle can obtain from its environment, driver and occupants, can be used to support personal assistants that give occupants the sensation of being in control — even if the vehicle is self-driving, potentially increasing not only the enjoyment of the vehicle but more importantly the perceived trustworthiness on the vehicle.
@@
Such assistants can be developed as a set of subscribing services within the vehicle’s ECUs dedicated to infotainment subsystems and, if using speech or touch interaction modalities, it can greatly benefit from a well defined timely delivery of information which decreases the overall jitter visible to the end users.
5.4 Services for assisted fleet management and maintenance
SDVs can also support scenarios related to management of vehicle(s), including managing its health, maintenance (predicatively or not) and homologation, both for individual vehicles or for a fleet.
5.4.1 Fleet management
Fleet management scenarios and remote services, such as the one presented at fig. 19626262More details at autowiz.in., are examples of use cases that can benefit from a set of services deployed across multiple vehicles. They consider a set of services containing the logic required to monitor, in real-time, the status (e.g. location or number of hours the vehicle is driven) of a vehicle within a fleet. This set of services, if orchestrated throughout a cluster of vehicles, can support systems for enterprises managing a vehicular fleet, homologating each vehicle, or even for state entities to perform the routine inspection. The information from/to those services, while it does not necessarily require an extremely low latency, it must be delivered and delivered within known time bounds and with low jitter.
5.4.2 Remote vehicle health and predictive maintenance
@@
Constant monitoring of the main vital parameters of a vehicle state — such as tire condition, brake system condition, engine cooling system status (in the case of internal combustion engines) or the power delivery subsystems (in the case of electric vehicles) — requires not only a reduced latency but also, in case the vehicle detects an issue, a guarantee of delivery in a bounded time. This is required so that, at the very least, other vehicles are aware of the problem before it has impact on the road. In the case of emergency scenarios — as previously described in section 5.2.2, the latency shall also be minimized.
@@
5.4.3 Over-the-air updates and real-time homologation
@@
Changes in the vehicles’ software — either via OTA updates, through the addition of new services or after a vehicle is sold to ensure no changes — should, in an automated way, be reflected in a re-homologation process. This process, while not requiring a minimized latency, shall be done in a bounded timely manner, with a guarantee of reliability; otherwise, vehicles may be driving on public roads with an unwanted, unsafe state [140].
6 Proposed vision
The SDV concept is capable of supporting several services spanning several different types of use cases. All the previously defined use cases require $N$ time critical services and $K$ services. In all these use cases, a certain vehicular service can either be fully deployed on one or more nodes, or divided through multiple nodes / cluster of vehicles, with both scenarios considering the cloud as an option. Also common to all use cases is their need for time criticality. Within the SDV paradigm, the vehicle may send the information about its occupant to other subsystems, or vehicles through communications, which must be deterministic in order to ensure the information is obtained not too soon nor too late, therefore risking lives. While the use cases do not necessarily focus on controlling the critical systems intra-vehicle such as power-train and braking subsystems, they are essential for other passive safety functionalities.
@@ 6 Proposed vision
Given the criticality on the one side of the use cases and associated services, and the increased complexity of development and deployment of services, it is essential to consider a dynamic management of services. Towards such management, several layers must be considered as depicted in fig. 20. These layers should act at least in the following domains [30]:
  • 1.
@@
    Orchestration of the services: where they should be run and policies to decide the node(s) and the number of replicas. Such domain is the responsibility of the orchestrator, a .
  • 2.
    Configuration of the abstraction layers required to support exchanging data in predefined formats and protocols. Such domain is the responsibility of the data layer configuration ( b ).
  • 3.
@@
    Management of the deterministic network: including automatic configuration of routing and deterministic traffic shapers and reliability processes. This domain is the responsibility of the deterministic network configurator ( c ).
  • 4.
@@
    Virtualization of the hardware in order to support better resource usage and decoupling of services and operating systems from each ECU. This domain is the responsibility of the vehicle abstraction layer ( d ).
@@
Figure 20: Overall layered view of the architecture main modules framed within the OSI model.
@@
Figure 21 presents the overview of an initial proposal of architecture towards such a dynamic management of vehicular services. This architecture supports support service scheduling, redundancy, and healing based on the vehicles’ computing capabilities.
@@
Figure 21: Overall high level architecture of the proposed solution, depicting the service development and deployment flow from the service developers down to the vehicular hardware.
@@
The Controller ( a ) is the main entry-point of the system to be researched and ensures the first aforementioned layer. Through user-defined and system-defined metrics within a Service Descriptor, it decides the amount of resources and the placement of those resources throughout the computing continuum and the different vehicles throughout the smart city, while considering possible handovers and optimal usage of the computing and network resources. The controller is responsible for controlling 4 verticals.
@@
The first vertical, the Deterministic Network ( b ), ensures that messages – either control messages to actuate or status messages containing information from vehicles and the smart city – are delivered in a well-specified time bound and with fault-proof reliability. Research topics here include, but are not limited to, integrating clock synchronization mechanisms into wireless vehicular networks and mechanisms towards dynamic automatic definition, and optimal configuration of flows of packets, their queue priorities and characteristics. The module within the architecture, the Deterministic Network Configurator, is responsible for creating the necessary traffic flows, their parameters and configuring the necessary hardware, namely the Network Interface Controller/Card (NIC) within each ECU and the TSN capable Ethernet switches interconnecting them. It is expected at least 3 traffic classes to be configured: a) system control messages with maximum criticality; b) high size (non) periodic streams of data and c) periodic or non periodic service messages.
@@
The Data Layer vertical ( c ) is responsible for ensuring the data is exchanged between the vehicular services in the most appropriated set of SOA protocols — such as SOME/IP, MQTT or DDS. Research topics here include, but are not limited to, the design of a abstraction layer. Such abstraction layer would increase interoperability between all these protocols while ensuring deterministic communication through integration with the underlying deterministic network mechanisms.
@@
The Orchestrator vertical ( d ) is responsible for ensuring the deployment of the services, either within a Kubernetes worker node, within a virtual machine managed for example by KVM or through a containerd or equivalent container runtime. Research topics here include ensuring deterministic communication through integration with the underlying deterministic network mechanisms (as previously mentioned in the Data Layer vertical).
@@
These 3 verticals dispatch orders to the Vehicle Abstraction Layer (VAL) ( e ) of each ECU, which abstracts different types and models of vehicles’ ECUs – both simulators and real vehicles – their actuators, sensors, and ECUs through the abstraction of existing frameworks, such as AUTOSAR, Autoware or others. Conceptually, the VAL is no different to the standardized abstractions operating systems created during the mid 1980s and 1990s to abstract the hardware from the remaining kernel, with [133] proposing an extension of such a paradigm into the vehicular paradigm, as depicted in fig. 22.
Figure 22: The VAL based architecture for vehicles,as described by [133], where the divisions between trusted and untrusted zones and between safety and non-safety services are evident
@@
6.1 Overall end-to-end interaction
Figure 23 provides an overview of the end-to-end interaction of the proposed system, covering everything from the features available to the end-user to the configuration of the required hardware needed to run the services and meet their requirements.